Please solve the linear differential equations
1. dx/y+z = dy/-(x+z) = dz/x-y

2.(x^2D^2-3xD+5)y =x^2(logx)
1
Expert's answer
2020-09-23T18:05:29-0400
"Solution"

1. dx/y+z = dy/-(x+z) = dz/x-y

We solve it y Lagrange's multipliers method


"\frac{(ldx+mdy+ndz)}{(lP+mQ+nR)}=0\\
lP+mQ+nR=0\\"

So,

"ldx+mdy+ndz=0"

Here "l,m,n" are multipliers

"\frac{dx}{y+z}=\frac{dy}{-x-z}=\frac{dz}{x-y}," given.


So,

"1dx+1dy+1dz=0"


Since

"1(y+z)+1(-x-z)+1(x-y)=0,"

"1,1,1" are multipliers

And

"x(y+z)+y(-x-z)+z(x-y)=0,"

So "x,y,z" are multipliers


"\therefore 1dx+1dy+1dz=0"

After integration

"x+y+z=C_1---------(i)"

Also "xdx+ydy+zdz=0"

After integration

"x^2+y^2+z^2=C_2---------(ii)"


"(i)" and "(ii)" are the required solutions


2.(x^2D^2-3xD+5)y =x^2(logx)


Let "x=e^z" and "ln x=z"

"\implies \frac{dx}{dz}=e^z=x\\
\implies \frac{dz}{dx}=\frac{1}{x}"


Now,

"\frac{dy}{dx}=\frac{dy}{dz} \cdot \frac{dz}{dx} = \frac{dy}{dz} \cdot \frac{1}{x}\\

x \frac{dy}{dx} = \frac{dy}{dz}\\
if\ \frac{d}{dz} \equiv \theta, then\\
xDy= \theta y"


Similarly,

"x^2D^2y=\theta(\theta - 1) y"

Now,

"(\theta ^2 - \theta - 3\theta+5)y=ze^{2z}\\
\implies (\theta ^2 - 4\theta+5)y=ze^{2z}"

This is a non homogeneous differential equation of the second order.


The remaining solution should be elementary.


The characteristic equation "r^2-4r+5" has the solutions "r=2-i, 2+i" . While the particular solution (by the method of undetermined coefficients) comes out to be "Y_p=e^{2z}"


Then the general solution (in terms of "z" ) is given by,

"y=y_c+y_p=C_1e^{z(2+i)}+C_2e^{z(2-i)}+e^{2z}"

for arbitrary constants "C_1" and "C_2."

"y=C_1x^{(2+i)}+C_2x^{(2-i)}+x^{2}"